Reports 1-1 of 1 Clear search Modify search
AdV-COM (AdV commissioning (1st part) )
mwas - 18:52 Saturday 14 February 2026 (68685) Print this report
Etalon heating belt power supply failing

Figure 1. The temperature of the NI etalon control started dropping today. Looking at the monitoring of the heating belt output it became intermitent.

This issue has appeared several times in the past and was solved by replacing the Kert power supply: https://logbook.virgo-gw.eu/virgo/?r=67582

We should consider starting to make the replacement of these power supplies on fixed schedule based on past expererience (for example changing them every year or two) instead of waiting for the power supply to fail and then have a thermal transient in the etalon loop lasting a for several days.

Images attached to this report
Comments to this report:
dattilo - 21:02 Sunday 15 February 2026 (68687) Print this report
This evening I went in TCS room where the power supply is located, and also looking at its analog display, it was possible to see that it was delivering intermittently, as indicated by the LSC_NI_HB_moni signal.
Therefore a around 7pm UTC I replaced the power supply with a spare one of the same kind (Kert 420). During the intervention I temporarily opened the NI etalon loop.
Now the loop seems working regularly (attached fig).
Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×